# A Training-Free Style-aligned Image Generation with Scale-wise Autoregressive Model

### v1

@@ Abstract
We present a training-free style-aligned image generation method that leverages a scale-wise autoregressive model. While large-scale text-to-image (T2I) models, particularly diffusion-based methods, have demonstrated impressive generation quality, they often suffer from style misalignment across generated image sets and slow inference speeds, limiting their practical usability. To address these issues, we propose three key components: initial feature replacement to ensure consistent background appearance, pivotal feature interpolation to align object placement, and dynamic style injection, which reinforces style consistency using a schedule function. Unlike previous methods requiring fine-tuning or additional training, our approach maintains fast inference while preserving individual content details. Extensive experiments show that our method achieves generation quality comparable to competing approaches, significantly improves style alignment, and delivers inference speeds over six times faster than the fastest model.
@@ 1 Introduction
Large-scale text-to-image (T2I) models [38, 37, 41, 17, 34, 5, 46] have become essential tools across various creative fields, including digital content creation, game design, advertising, and artistic visualization.
However, the growing use of these models has revealed a key limitation: style misalignment, where images fail to maintain a consistent visual style across objects, prompting various research efforts [43, 13, 19, 56, 39]. One prominent approach leverages parameter-efficient fine-tuning (PEFT), such as Low-Rank Adaptation (LoRA) [22], which has been applied to specific components of diffusion architectures [43, 13]. These methods enable efficient style alignment with minimal computational overhead during training. While effective, these methods still require costly additional fine-tuning. Alternatively, training-free methods have also been proposed [19], such as replacing the self-attention layer with a shared attention layer to enforce style consistency. Despite these efforts, the inherent issue of long inference times in diffusion-based models remains unresolved.
As a faster alternative, vector quantized (VQ)-based autoregressive models [49, 12] have been proposed to generate images by predicting discrete tokens. The emergence of mask schedule-based non-autoregressive transformers [4, 27, 5], inspired by BERT [10], has further improved both the generation speed and quality of these models. Building on this progress, the recently proposed VAR-based next-scale prediction method [47] has pushed the performance of autoregressive models even closer to that of diffusion models [46, 17], while maintaining significantly faster inference times. However, similar to diffusion-based models, autoregressive models also struggle with a style misalignment issue, as illustrated in Fig. 1. LoRA-based style fine-tuning has been applied to non-autoregressive transformers [44] to address this issue, but it remains computationally expensive and time-consuming due to the additional training required.
@@
  • To the best of our knowledge, this is the first work to propose a training-free style-aligned image generation based on an autoregressive model, achieving over 6$\times$ faster inference along with the best style alignment performance.
  • We introduce initial feature replacement and pivotal feature interpolation to ensure consistent RGB statistics and object placement across generated images.
  • We present dynamic style injection, which enhances style consistency while preserving content fidelity.
@@ 2 Related works
Figure 2: Inference time ($\downarrow$, lower is better) vs. dual consistency ($\uparrow$, higher is better) curve comparing ours with competitive methods (StyleAligned [19], B-LoRA [13], StyleDrop [44], DreamBooth-LoRA (DB-LoRA) [40], and IP-Adapter [56]).
Figure 3: Visualization of images generated at different steps of the next-scale prediction process. In the early and mid-stages, global composition and overall style are established, while later steps focus on refining details and textures. We also track RGB statistics, content similarity, and style similarity across 400 generated images, comparing each step to the final output (12th step) to evaluate the progression of content preservation, style consistency, and RGB statistics.
@@ 2.1 Text-to-image generation
The development of large-scale text-image pair datasets [42, 2, 6, 32], combined with advances in generative models such as diffusion models [20, 45], GANs [16], and autoregressive models [49, 12], has significantly accelerated the progress of large-scale text-to-image (T2I) generation models [38, 37, 41, 17, 34, 5, 46, 26]. In particular, diffusion-based text-to-image models have demonstrated outstanding performance, making them widely adopted for various downstream applications such as style transfer and image editing [9, 54, 48, 1, 25, 18, 35]. More recently, autoregressive models employing next-scale prediction [47] have emerged as promising alternatives, offering significantly faster inference speeds while maintaining competitive generation quality compared to diffusion-based approaches. This advancement has opened new avenues for efficient T2I generation [17, 46]. Despite these advancements, both diffusion and autoregressive models still suffer from persistent challenges in style alignment, which limits their practical usability and degrades user experience.
@@ 2.2 Style Transfer
Image style transfer, pioneered by [14], leverages pre-trained CNNs like VGGNet to extract content and style features, laying the groundwork for this research area. However, these early optimization-based methods suffer from high computational costs, as they require per-image optimization. To mitigate this, Huang et al. [24] introduces Adaptive Instance Normalization (AdaIN) for real-time style transfer, followed by Whitening and Coloring Transform (WCT) [30, 31], which improves feature alignment. With the rise of attention mechanisms [50, 11], newer models [33, 21, 55] have further enhanced stylization quality. Generative models such as GANs [16], diffusion models [20], and vector quantization-based autoregressive models [12] have opened new possibilities in style transfer. Diffusion-based methods [9, 54, 28] achieve high visual fidelity by integrating style features through cross-attention during the denoising process, while vector quantization approaches [7, 53, 23, 15] enhance stylization by combining content tokenization with learned style representations. Despite their advantages, early vector quantization models often require separate style codebooks or extensive fine-tuning, while diffusion-based methods face challenges due to their long inference times.
@@ 2.3 Personalized image generation
Various personalized image generation methods have been proposed to adapt new visual features to user intent using pre-trained text-to-image models. These methods can be broadly categorized into content-oriented and style-oriented methods. Content-oriented methods [39, 29, 52] primarily focus on capturing object-specific attributes or synthetic features to generate images that explicitly reflect a target subject. They leverage pre-trained models or fine-tuning methods to embed subject-specific characteristics into the generated images. In contrast, style-oriented methods [44, 13, 43, 56, 40] focus on controlling the style of the generated images. In particular, Style-Aligned Generation [19], which closely relates to our approach, enhances style-consistency across batches by sharing attention during the denoising diffusion process and reducing differences among generated images through AdaIN [24].
@@ 3 Method
[…]
@@ 3.1 Observation of next-scale prediction
To gain deeper insight into the internal mechanisms underlying the next-scale prediction scheme, we analyze the evolution of RGB statistics, style, and content similarity throughout the autoregressive generation process, as illustrated in Fig. 3. To assess the progression of content preservation, style consistency, and RGB statistics over time, we compare the outputs at each step to the final output (the 12th step) using 400 generated images. RGB statistics are quantified using the chi-square distance between histograms. Content similarity is measured using the cosine similarity of VGG19 features, while style similarity is evaluated using the average pairwise cosine similarity of DINO ViT-B/8 [3] embeddings within each generated set. Based on this analysis, we identify the following key observations, which form the foundation of our style-aligned image generation method.
@@
(1) Dominance of RGB statistics in early-stage: As shown in Fig. 3, the early-stage (e.g., the 1st-2nd steps) primarily establish the overall RGB statistics of the generated images. Once the RGB statistics are set, the autoregressive model, being heavily conditioned on the outputs from previous steps, follows these statistics closely, strongly influencing all subsequent stages. This means that any modifications made during these early-stage have a significant impact on the final RGB statistics. This observation highlights the importance of controlling early-stage representations to ensure consistent style across a set of generated images. Based on this observation, we propose the initial feature replacement method described in Sec. 3.3 enforcing consistent RGB statistics during the early-stage to promote style alignment across generated images.
@@
(2) Dynamic transition of style and content in mid-stage: Fig. 3 further reveals that both style and content gradually evolve with increasing detail during the mid-stage, eventually stabilizing as the generation process approaches the later steps (e.g., the 8th step). Due to the autoregressive nature of the model, early-stage decisions strongly influence the overall generation flow, shaping both style and content composition throughout the process. This observation aligns with findings from [51], demonstrating that modifications in the late-stage have minimal impact on the overall semantic structure, while adjustments made during the mid-stage effectively blend both style and content attributes. Based on this, we identify the mid-stage as an optimal window for style refinement enabling us to enhance style consistency without significantly altering the underlying content structure. This insight serves as the foundation of our strategy to apply targeted interventions during these intermediate steps, harmonizing style and content before they fully stabilize in the late-stage.
@@
To leverage this observation, we propose the pivotal feature interpolation described in Sec. 3.4, which aligns the object’s silhouette and ensures uniform style by adjusting features at one of the mid-stage. In addition, we introduce the style injection in Sec. 3.5, which dynamically injects style using a designed scheduling function. This ensures that style consistency is continuously reinforced throughout the generation process, while allowing content details to be progressively refined without distortion.
@@ 3.2 Overall pipeline
In this paper, we aim to generate a set of images $\mathbf{I}=\{I^{n}\}_{n=1}^{N}$ corresponding to an input set of text prompts $\mathbf{T}=\{T^{n}\}_{n=1}^{N}$, while ensuring that all generated images share a consistent visual style. To achieve this, we concatenate the text prompts and process them in parallel as a batch, enabling the model to generate multiple images simultaneously with shared stylistic coherence. The overall pipeline of our model is illustrated in Fig. 4. Our approach builds upon the architecture of Infinity [17], a state-of-the-art T2I model based on the next-scale prediction paradigm [47]. Specifically, our model consists of a pre-trained text encoder $E_{T}$ derived from Flan-T5 [8], an autoregressive transformer model $M$ responsible for next-scale prediction, and a decoder $D$ that reconstructs the final image from the predicted residual feature maps. Based on observation in Sec. 3.1, we define the set of all generation steps as $\mathbf{S}=\{1,2,\dots,S\}$, the early steps as $\mathbf{S}_{e}=\{1,2\}$, and middle steps as $\mathbf{S}_{m}=\{3,\dots,7\}$.
[…]
@@
$\mathbf{F}_{s}=\sum_{i=1}^{s}\text{up}_{H\times W}(\mathbf{R}_{i}),~{}~{}\mathbf{R}_{s}\in\mathbb{R}^{N\times h_{s}\times w_{s}}\$ (2)
@@ 3.3 Initial feature replacement
As described in Sec. 3.1, the initial steps in next-scale prediction play a dominant role in establishing the overall RGB statistics of the generated images, while contributing minimally to fine-grained content. This observation aligns closely with the core design philosophy of next-scale prediction in VAR [47], where the multi-scale, coarse-to-fine process naturally enforces a structured “order” on image generation.
@@
$\begin{gathered}\mathbf{F}_{s}\leftarrow\{\tilde{F}_{s}^{n}\}_{n=1}^{N},~{}\forall{s}\in\mathbf{S}_{\text{e}},\\ \tilde{F}_{1}^{n}=F_{1}^{1},~{}\forall n\in\mathbf{N},\\ \end{gathered}$ (4)
where $\mathbf{N}=\{1,2,\dots,N\}$ denotes the set of image indices. This replacement helps ensure that the images generated in later steps within a batch share consistent and harmonious color statistics, while still preserving the unique content and identity of each individual image.
@@ 3.4 Pivotal feature interpolation
The style-aligned text-to-image generation task [19] aims to produce a set of images that maintain consistent object placement and a uniform style across all generated images. Based on our observations in Sec. 3.1, we identify that one of the mid-stage steps in the coarse-to-fine next-scale prediction process plays a key role in determining both object placement and overall style. To leverage this property, we apply feature interpolation at $\bar{s}$-th step within the mid-stage $\mathbf{S}_{\text{m}}$, where each feature is interpolated with the first feature in the batch. The feature interpolation is defined as:
$\begin{gathered}\mathbf{F}_{\bar{s}}\leftarrow\{\bar{F}_{\bar{s}}^{n}\}_{n=1}^{N},~{}\bar{s}\in\mathbf{S}_{\text{m}},\\ ~{}\bar{F}_{\bar{s}}^{n}=\alpha_{\text{pivot}}{F}_{\bar{s}}^{1}+(1-\alpha_{\text{pivot}})\bar{F}_{\bar{s}}^{n},~{}\forall n\in\mathbf{N},\end{gathered}$ (5)
where $\alpha_{\text{pivot}}$ denotes interpolation weight. This guides the generation process ensuring consistent object placement and visual style across the generated images.

### v2

@@ Abstract
[…]
@@ 1 Introduction
Large-scale text-to-image (T2I) models [rombach2022highresolution, ramesh2021zeroshot, saharia2022photorealistictexttoimagediffusionmodels, han2024infinity, podell2023sdxlimprovinglatentdiffusion, chang2023musetexttoimagegenerationmasked, tang2024hartefficientvisualgeneration] have become essential tools across various creative fields, including digital content creation, game design, advertising, and artistic visualization.
However, the growing use of these models has revealed a key limitation: style misalignment, where images fail to maintain a consistent visual style across objects, prompting various research efforts [shah2023ziplorasubjectstyleeffectively, frenkel2024implicitstylecontentseparationusing, hertz2024stylealignedimagegeneration, ye2023ipadaptertextcompatibleimage, ruiz2023dreambooth]. One prominent approach leverages parameter-efficient fine-tuning (PEFT), such as Low-Rank Adaptation (LoRA) [hu2021loralowrankadaptationlarge], which has been applied to specific components of diffusion architectures [shah2023ziplorasubjectstyleeffectively, frenkel2024implicitstylecontentseparationusing]. These methods enable efficient style alignment with minimal computational overhead during training. While effective, these methods still require costly additional fine-tuning. Alternatively, training-free methods have also been proposed [hertz2024stylealignedimagegeneration, zhang2025alignedgen], such as modifying the self-attention layer with a shared attention layer to enforce style consistency. Despite these efforts, the inherent issue of long inference times in diffusion-based models remains unresolved.
As a faster alternative, vector quantized (VQ)-based autoregressive models [van2017neural, esser2021taming] have been proposed to generate images by predicting discrete tokens. The emergence of mask schedule-based non-autoregressive transformers [chang2022maskgit, kondratyuk2023videopoet, chang2023musetexttoimagegenerationmasked], inspired by BERT [devlin2019bert], has further improved both the generation speed and quality of these models. Building on this progress, the recently proposed VAR-based next-scale prediction method [tian2024visual] has pushed the performance of autoregressive models even closer to that of diffusion models [tang2024hartefficientvisualgeneration, han2024infinity], while maintaining significantly faster inference times. However, similar to diffusion-based models, autoregressive models also struggle with a style misalignment issue, as illustrated in Fig. 1. LoRA-based style fine-tuning has been applied to non-autoregressive transformers [sohn2023styledrop] to address this issue, but it remains computationally expensive and time-consuming due to the additional training required.
@@
  • We conduct a comprehensive analysis of the scale-wise autoregressive model, revealing the distinct roles of each generation step.
  • Based on our analysis, we introduce three simple yet effective techniques—initial feature replacement, pivotal feature interpolation, and dynamic style injection—to enhance style consistency while preserving content fidelity.
[…]
@@ 2 Related works
Figure 2: Inference time ($\downarrow$, lower is better) vs. dual consistency ($\uparrow$, higher is better) curve comparing ours with competitive methods (StyleAligned [hertz2024stylealignedimagegeneration], B-LoRA [frenkel2024implicitstylecontentseparationusing], StyleDrop [sohn2023styledrop], DreamBooth-LoRA (DB-LoRA) [ryu2023low], IP-Adapter [ye2023ipadaptertextcompatibleimage], AlignedGen [zhang2025alignedgen], CSGO [xing2024csgo] and DreamO [mou2025dreamo]).
Figure 3: Visualization of images generated at different steps of the next-scale prediction process. In the early- and mid-stages, global composition and overall style are established, while later steps focus on refining details and textures. We also track RGB statistics, content similarity, and style similarity across 400 generated images, comparing each step to the final output (12th step) to evaluate the progression of content preservation, style consistency, and RGB statistics.
@@ 2.1 Text-to-image generation
The development of large-scale text-image pair datasets [schuhmann2022laion5bopenlargescaledataset, kakaobrain2022coyo-700m, changpinyo2021conceptual12mpushingwebscale, lin2015microsoft], combined with advances in generative models such as diffusion models [ho2020denoising, song2022denoising], GANs [goodfellow2014generative], and autoregressive models [van2017neural, esser2021taming], has significantly accelerated the progress of large-scale text-to-image (T2I) generation models [rombach2022highresolution, ramesh2021zeroshot, saharia2022photorealistictexttoimagediffusionmodels, han2024infinity, podell2023sdxlimprovinglatentdiffusion, chang2023musetexttoimagegenerationmasked, tang2024hartefficientvisualgeneration, kang2023gigagan]. In particular, diffusion-based text-to-image models have demonstrated outstanding performance, making them widely adopted for various downstream applications such as style transfer and image editing [chung2024style, tumanyan2022plugandplay, brooks2023instructpix2pix, jeong2024trainingfree]. More recently, autoregressive models employing next-scale prediction [tian2024visual] have emerged as promising alternatives, offering significantly faster inference speeds while maintaining competitive generation quality compared to diffusion-based approaches. This advancement has opened new avenues for efficient T2I generation [han2024infinity, tang2024hartefficientvisualgeneration]. Despite these advancements, both diffusion and autoregressive models still suffer from persistent challenges in style alignment, which limits their practical usability and degrades user experience.
@@ 2.2 Style Transfer
Image style transfer, pioneered by [Gatys_2016_CVPR], leverages pre-trained CNNs like VGGNet to extract content and style features, laying the groundwork for this research area. However, these early optimization-based methods suffer from high computational costs, as they require per-image optimization. To mitigate this, Huang et al. [Huang_2017_ICCV] introduces Adaptive Instance Normalization (AdaIN) for real-time style transfer, followed by Whitening and Coloring Transform (WCT) [li2017universal, li2018closed], which improves feature alignment. With the rise of attention mechanisms [vaswani2023attention, dosovitskiy2021image], newer models [liu2021adaattn, hong2023aespanet, yao2019attentionaware] have further enhanced stylization quality. Generative models such as GANs [goodfellow2014generative], diffusion models [ho2020denoising], and vector quantization-based autoregressive models [esser2021taming] have opened new possibilities in style transfer. Diffusion-based methods [chung2024style, yang2023zero, kwon2023diffusionbased] achieve high visual fidelity by integrating style features through cross-attention during the denoising process, while vector quantization approaches [chen2022vector, xu2023stylerdalle, huang2023quantart, Gim_2024_ACCV] enhance stylization by combining content tokenization with learned style representations. Despite their advantages, early vector quantization models often require separate style codebooks or extensive fine-tuning, while diffusion-based methods face challenges due to their long inference times.
@@ 2.3 Personalized image generation
Various personalized image generation methods have been proposed to adapt new visual features to user intent using pre-trained text-to-image models. These methods can be broadly categorized into content-oriented and style-oriented methods. Content-oriented methods [ruiz2023dreambooth, li2023blip, wei2023elite] primarily focus on capturing object-specific attributes or synthetic features to generate images that explicitly reflect a target subject. They leverage pre-trained models or fine-tuning methods to embed subject-specific characteristics into the generated images. In contrast, style-oriented methods [sohn2023styledrop, frenkel2024implicitstylecontentseparationusing, ryu2023low, zhang2025alignedgen, xing2024csgo, mou2025dreamo] focus on controlling the style of the generated images. In particular, Style-Aligned Generation [hertz2024stylealignedimagegeneration], which closely relates to our approach, enhances style-consistency across batches by sharing attention during the denoising diffusion process and reducing differences among generated images through AdaIN.
@@ 3 Method
Figure 4: Overall pipeline of our model. The T5 text encoder processes text prompts $\mathbf{T}$, providing conditions and $\langle\text{SOS}\rangle$ tokens to the transformer. Initial Feature Replacement aligns RGB statistics at the 1st and 2nd generation steps. Pivotal Feature Interpolation adjusts object positions and styles at the $\bar{s}$-th step ($\bar{s}$=3), while Dynamic Style Injection gradually reduces style influence from the 3rd to 7th steps. The decoder converts the transformer’s final step output $\mathbf{F}_{S}$ into the style-aligned images $\mathbf{I}$.
@@ 3.1 Observation of next-scale prediction
To gain deeper insight into the internal mechanisms underlying the next-scale prediction scheme, we analyze the evolution of RGB statistics, style, and content similarity throughout the autoregressive generation process, as illustrated in Fig. 3. We evaluate the progression of content preservation, style consistency, and RGB statistics by comparing each generation step to the final output (the 12th step) using 400 images generated from randomly sampled text prompts created with ChatGPT. RGB statistics are quantified using the chi-square distance between histograms, content similarity with the cosine similarity of VGG19 features [simonyan2014very], and style similarity with the average pairwise cosine similarity of DINO ViT-B/8 [caron2021emerging] embeddings within each set. Based on this analysis, we identify the following key observations, which form the foundation of our method.
@@
(1) Dominance of RGB statistics in early-stage: As shown in Fig. 3, the early-stage (e.g., the 1st-2nd steps) primarily establishes the overall RGB statistics of the generated images. Once the RGB statistics are set, the autoregressive model, heavily conditioned on the outputs from previous steps, closely follows these statistics, strongly influencing all subsequent stages. This means that any modifications made during the early-stage have a significant impact on the final RGB statistics. This observation highlights the importance of controlling early-stage representations to ensure consistent style across a set of generated images. Based on this observation, we propose the initial feature replacement method described in Sec. 3.3, enforcing consistent RGB statistics during the early-stage to promote style alignment across generated images.
@@
(2) Dynamic transition of style and content in mid-stage: Fig. 3 further reveals that both style and content gradually evolve with increasing detail during the mid-stage, eventually stabilizing as the generation process approaches the later steps (e.g., the 8th step). Due to the autoregressive nature of the model, early-stage decisions strongly influence the overall generation flow, shaping both style and content composition throughout the process. This observation aligns with findings from [voronov2024swittidesigningscalewisetransformers], demonstrating that modifications in the late-stage have minimal impact on the overall semantic structure, while adjustments made during the mid-stage effectively blend both style and content attributes. Based on this, we identify the mid-stage as an optimal window for style refinement, enabling us to enhance style consistency without significantly altering the underlying content structure. This insight serves as the foundation of our strategy to apply targeted interventions during these intermediate steps, harmonizing style and content before they fully stabilize in the late-stage.
(3) Quantitative analysis of feature dynamics: To further substantiate these observations, we conduct a quantitative analysis of the scale-wise generation process, focusing on feature variations across early, mid, and late-stages. Specifically, we measure the average differences between consecutive generation steps in both content and style features. For each region of the generation process—early (steps 1–2), mid (steps 3–7), and late (steps 8–12)—we compute the mean feature changes as follows:
  • Content feature differences: Early: 0.031, Mid: 0.137, Late: 0.042
  • Style feature differences: Early: 0.002, Mid: 0.158, Late: 0.018
These statistics quantitatively confirm that both content and style features vary most significantly during the mid-stage, validating our stage-wise decomposition of the generation process into early, mid, and late phases.
@@
To leverage this observation, we propose the pivotal feature interpolation described in Sec. 3.4, which aligns the object’s silhouette and ensures uniform style by adjusting features at one of the steps in mid-stage. In addition, we introduce the dynamic style injection in Sec. 3.5, which dynamically injects style using a designed scheduling function. This ensures that style consistency is continuously reinforced throughout the generation process, while allowing content details to be progressively refined without distortion.
@@ 3.2 Overall pipeline
In this paper, we aim to generate a set of images $\mathbf{I}=\{I^{n}\}_{n=1}^{N}$ corresponding to an input set of text prompts $\mathbf{T}=\{T^{n}\}_{n=1}^{N}$, while ensuring that all generated images share a consistent visual style. To achieve this, we concatenate the text prompts and process them in parallel as a batch, enabling the model to generate multiple images simultaneously with shared stylistic coherence. The overall pipeline of our model is illustrated in Fig. 4. Our approach builds upon the architecture of Infinity [han2024infinity], a state-of-the-art T2I model based on the next-scale prediction paradigm [tian2024visual]. Specifically, our model consists of a pre-trained text encoder $E_{T}$ derived from Flan-T5 [chung2022scalinginstructionfinetunedlanguagemodels], an autoregressive transformer model $M$ responsible for next-scale prediction, and a decoder $D$ that reconstructs the final image from the predicted residual feature maps. Based on our observations, we define the set of all generation steps as $\mathbf{S}=\{1,2,\dots,S\}$, the early steps as $\mathbf{S}_{\text{e}}=\{1,2\}$, and middle steps as $\mathbf{S}_{\text{m}}=\{3,\dots,7\}$.
The autoregressive model $M$ predicts a sequence of quantized residual feature maps $\mathbf{R}_{s}=\{R^{n}_{s}\}_{n=1}^{N}$, conditioned on the input text prompts $\mathbf{T}$ and previously generated features $\mathbf{F}_{s-1}$. The initial features $\mathbf{F}_{0}$ correspond to the start-of-sequence $\langle\text{SOS}\rangle$ tokens. The prediction process at all generation steps $s\in\mathbf{S}$ can be described as follows:
@@
$\mathbf{F}_{s}=\sum_{i=1}^{s}\text{up}_{H\times W}(\mathbf{R}_{i}),~~\mathbf{R}_{s}\in\mathbb{R}^{N\times h_{s}\times w_{s}}\\$ (2)
@@ 3.3 Initial feature replacement
As described in Sec. 3.1, the initial steps in next-scale prediction play a dominant role in establishing the overall RGB statistics of the generated images, while contributing minimally to fine-grained content. This observation aligns closely with the core design philosophy of next-scale prediction in VAR [tian2024visual], where the multi-scale, coarse-to-fine process naturally enforces a structured “order” on image generation.
@@
[…]
where $\mathbf{N}=\{1,2,\dots,N\}$ denotes the set of image indices. This replacement helps ensure that the images generated in later steps within a batch share consistent and harmonious RGB statistics, while still preserving the unique content and identity of each individual image.
@@ 3.4 Pivotal feature interpolation
The style-aligned text-to-image generation task [hertz2024stylealignedimagegeneration] aims to produce a set of images that maintain consistent object placement and a uniform style across all generated images. Based on our observations in Sec. 3.1, we identify that one of the mid-stage steps in the coarse-to-fine next-scale prediction process plays a key role in determining both object placement and overall style. To leverage this property, we apply feature interpolation at $\bar{s}$-th step within the mid-stage $\mathbf{S}_{\text{m}}$, where each feature is interpolated with the first feature in the batch. The feature interpolation is defined as:
$\begin{gathered}\mathbf{F}_{\bar{s}}\leftarrow\{\bar{F}_{\bar{s}}^{n}\}_{n=1}^{N},~\bar{s}\in\mathbf{S}_{\text{m}},\\ ~\bar{F}_{\bar{s}}^{n}=\alpha_{\text{pivot}}{F}_{\bar{s}}^{1}+(1-\alpha_{\text{pivot}}){F}_{\bar{s}}^{n},~\forall n\in\mathbf{N},\end{gathered}$ (5)
where $\alpha_{\text{pivot}}$ denotes interpolation weight. This guides the generation process, ensuring consistent object placement and visual style across the generated images.